OpenStudy (anonymous):

Help Please Please Please! (: Optimization: A rectangle has one side on the x-axis and two vertices on the curve y=3 / x^2+2. Find the vertices of the rectangle with maximum area.

OpenStudy (anonymous):



OpenStudy (anonymous):

not nearly as hard as it looks base will be \(2x\) and height will be \(\frac{3}{x^2+2}\) therefore the area of the rectangle will be \[A(x)=\frac{6x}{x^2+2}\] maximize that one

OpenStudy (anonymous):

I find A'(x) right? @satellite73

OpenStudy (anonymous):

\[A'(x)= \frac{ 12x }{ x^2+2 } - \frac{ 12x^3 }{ (x^2+2)^2 }\]

OpenStudy (anonymous):

\[A'(x)=0\] x=0 @satellite73

OpenStudy (mathstudent55):

I got \( A'(x) = \dfrac{-12x^2} {(x^2 + 2)^2} + \dfrac{6}{x^2 + 2} \)

OpenStudy (mathstudent55):

It can be simplified to: \(A'(x) = \dfrac{-6x^2 + 12}{(x^2 + 2)^2} \)

OpenStudy (anonymous):

how'd you get a negative 12?

OpenStudy (anonymous):

nevermind I caught my mistake.. @mathstudent55

OpenStudy (anonymous):

\[x=\pm \sqrt{2}\]

OpenStudy (anonymous):

right? @mathstudent55

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

So where so I go from there? @mathstudent55

OpenStudy (mathstudent55):

Now plug in those two values of x in the original function to find the corresponding y-coordinates of the two points on the function.

OpenStudy (anonymous):

you are done

OpenStudy (anonymous):

the original finction of y= or a(x)=?

OpenStudy (anonymous):

\(A(x)\) will give you the maximum area, but you are asked for the vertices

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

so they are \((-\sqrt2,0)\) and \((\sqrt2,0)\) for the ones on the \(x\) axis

OpenStudy (mathstudent55):

The function.

OpenStudy (anonymous):